OpenStudy (anonymous):

A baseball bat originally sold for $65 is on sale for $50. what percent off is the baseball bat? having trouble doing percentages

OpenStudy (anonymous):

Are there answer choices?

OpenStudy (anonymous):

there is A. 0.13% B 13% C 77% D. 0.77%

OpenStudy (anonymous):

\(\sf \Large\frac{new\ price}{original\ price} \times 100\)

OpenStudy (anonymous):

divide the new amount by the original, so 50/65. This number times 100 gives you the percent amount that the new number is of the old one. In essence, 50 is blank percent of 65. However, you don't want that number. You want the percent off, which would be the remaining percent. So subtract the number that you get from 50/65 times 100 and that's your answer.

OpenStudy (anonymous):

that's what I posted before but the answer calculated is not one of the choices

OpenStudy (anonymous):

50/65=.77 100 x .77=77 100-77=13 so B

OpenStudy (anonymous):

100-77=23 , not 13

OpenStudy (anonymous):

excuse me my bad thanks. Yeah 23 would be the answer.

OpenStudy (anonymous):

:) but it's not one of the choices so idk

OpenStudy (anonymous):

thank you very much this math is rediculious

OpenStudy (anonymous):

b